23.14 An ac generator has a frequency of 2.3 kHz and a voltage of 240 V. An inductance L1 = 4.0 mH is connected across its terminals. Then a second inductance L2 = 9.0 mH is connected in parallel with L1. Find (a) the current that the generator delivers to L1 and (b)to the parallel combination.
Answer #1

a] reactance X = 2pi f L1 = 2pi*2.3e3*4e-3 = 57.8 ohm

Current I = V/X = 240/57.8 = 4.152 A

b] Equivalent inductance = L1*L2/[L1+L2]

= 9*4/[9+4] = 2.769 mH

reactance X = 2pi f L = 2pi*2.3e3*2.769e-3 = 40.0157 ohm

Current I = V/X = 240/40.0157 = 6 A